Eric has 5 more dimes than quarters. He has 39 coins in all. How many dimes does Eric have? Identify the variable, write an equation and solve.
Answer by Boreal(15235)

quarters=x
dimes=x+5
x+x+5=39. x is the variable and is the number of dimes. But the number of quarters is related to the number of dimes in that there are 5 more dimes than quarters and their sum is 39.
2x+5=39
2x=34
x=17 quarters
x+5=22 dimes